OpenStudy (anonymous):

A private plane leaves an airport and flies due east at 192 miles per hour. Two hours later, a jet leaves the same airport and flies due east at 960 miles per hour. How far will the private plane fly before it is passed by the jet?

OpenStudy (anonymous):

@AnswerMyQuestions @helper99

OpenStudy (anonymous):

Let's say X is the amount of time that the first plane flies, and Y is the amount of time the second plane flies. X and Y are both in hours. We are looking for the point when the planes meet, so the distance that each plane flies will be the same. Distance is found by multiplying speed by time, so: 192x=960y Y is 2 hours less than X, so: x-2=y , so: 192x=960(x-2) 192x=960x-1920 768x=1920 x=2.5 So the first plane was flying for 2.5 hours before the second plane passed it.

OpenStudy (anonymous):

so 193x=______

OpenStudy (anonymous):

yes?